# Modelling two-dimensional droplet rebound off deep fluid baths

### v2

@@ Abstract
In order for a droplet to rebound rather than coalesce with a liquid bath, a layer of gas must persist throughout the impact. This gas, typically an air layer acts as a lubricant to the system and permits a pressure transfer between the two liquid bodies. Through considering separately the bath, air, and drop regions of fluid, we introduce a fully coupled reduced dynamic model of two-dimensional droplets rebounding off liquid baths, which incorporates an evolving lubricating air layer. Numerical solutions of the lubrication-mediated model are compared to dedicated direct numerical simulation (DNS) of the Navier-Stokes equations. The reduced model captures rebounding dynamics well in the regime where it is most relevant: for low-speed impacts of small droplets, where capillary forces are important. Numerically, the reduced model is efficient, allowing for the computation of multiple rebounds and of long time dynamics of droplets rebounding on a vibrating bath. Furthermore, the lubrication-mediated model is able to provide detailed information within the air layer such as pressure and lubrication-layer geometry, which is usually omitted from reduced models.
keywords:
@@ 1 Background and Motivation
Since Worthington (1881), we have endeavoured to classify the different behaviours that arise when droplets impact a liquid surface. An important question is to distinguish under which conditions a droplet will rebound versus coalesce or splash (Rein, 1993), especially in industrial processes such as spray injections (Moreira et al., 2010).
@@
The ability for a droplet to rebound depends entirely on whether the layer of air that separates the two fluids remains unbroken throughout the impact. In the event that this air layer becomes sufficiently thin, the Van der Waals forces between the two liquid regions act, and the liquid pierces through the air layer and initiates coalescence (Sprittles, 2024). For rebound to occur, it is a delicate balance between the system having enough restoring force to jettison the drop upwards after a cushioned impact, but not so much that the air layer is breached due to the inertia of the drop before the liquid surface has time to force the droplet upwards.
@@
Industrial interest in solid impact dynamics were initially heavily motivated by military interest in sea-plane landings and slamming of boat hulls with Wagner theory (Wagner, 1932; Korobkin, 2004). The role of the air layer in pre-impact dynamics of solid bodies on deep liquid baths was first developed by Smith et al. (2003). For this work we base ourselves on the investigation of Moore (2021), who combined pre-impact cushioning with the Wagner model. We incorporate surface tension effects in drop oscillations and in wave generation and propagation in the bath, relevant to scenarios such as spray coating and ink-jet printing (see Derby (2010)) where surface tension dominates small droplet behaviour.. Additionally, studies of droplet-droplet impacts have been motivated by the rich field of aerosols in a wide variety of natural processes from cloud seeding (Bird et al., 2010) to pathogen transfer through suspended droplets (Yang et al., 2007).
@@
This study is originally motivated by phenomena where a vertical vibration of the liquid bath permits sustained repeated droplet bouncing to occur Couder et al. (2005). These Faraday pilot-wave systems, named due to the Faraday waves generated at sufficiently strong vibration, can present a wide range of complex behaviours. To obtain a strongly coupled drop-bath system, the vibration is held below that which destabilises the free surface (the Faraday threshold). Within this regime, there are multiple drop-bath behaviours that can be experimentally achieved and modelled, and have been quantified in work such as Protiere et al. (2006), and Milewski et al. (2015). As the forcing reaches near the Faraday threshold, the bouncing modes are destabilised, and through the asymmetry of impact a horizontal force is applied to the drop, permitting them to ‘walk’ along the surface of the bath, guided by the waves generated from previous bounces. This Faraday pilot-wave system has drawn parallels to the quantum theory described by de Broglie (1926). Dubbed the hydrodynamic quantum analogue, attempts have been made to recreate macro-scale analogies to quantum phenomena. For comprehensive reviews see Bush (2015) and Bush et al. (2018).
@@
The linear inviscid model for the oscillations of a droplet was developed in Lamb (1924). The model has been validated experimentally by work such as in Becker et al. (1991). Considering droplets in zero gravity cases allows for the uninterrupted study of their oscillatory nature Wang et al. (1996) and the longer viscous damping timescales. Comparisons between computations of the Navier-Stokes equations and linear theory in two- and three-dimensions can be found in Aalilija et al. (2020). Following Lamb, they used the inviscid solution and an energy argument to find linear damping rates. For bouncing droplets, Moláček & Bush (2013) considered a simplified model where both bath and droplet deform equally as a logarithmic spring. Terwagne et al. (2013) experimentally investigated the role of deformation in the droplet rebound dynamics when the drop and bath are in different regimes, and proposed a double mass-spring system to replicate this range of regimes in a drop-bath system. While solid spheres can rebound in this regime Galeano-Rios et al. (2021), droplet deformation converts kinetic energy to drop deformations on contact with the bath, and then may revert it back to kinetic energy at take-off, increasing upwards propulsion.
@@
where here the Weber number indicates ratio of inertia to surface tension, the Reynolds number encapsulates the relative effect of inertia to viscosity, and the Ohnesorge number provides a relationship between viscosity and surface tension. There are several other parameters which could play a role; in particular for larger drops and longer waves the Bond number $Bo=\rho gR_{0}^{2}/\sigma$ gains importance. The work presented within this paper will focus around the low-Weber and the quasi-inviscid low-Ohnesorge limit, in order to capture smooth rebound dynamics with small disturbances (Sprittles, 2024).
The investigation of single rebounds in this regime has thrived over recent years, with the main goal being computationally efficient models that faithfully reproduce the experimental and DNS results. The ‘kinematic match’ method for axisymmetric (Galeano-Rios et al., 2017) and non-axisymmetric impacts (Galeano-Rios et al., 2019) modeled the droplet as a sphere, with the pressure on its contact surface obtained as a constraint for matching the shape of the sphere’s and the surface of the bath. This method captured details of bouncing and walking droplets, matching experiments, and was later used in comparisons of solid sphere impacts (Galeano-Rios et al., 2021). The model neglected droplet deformation and the air layer. Alventosa et al. (2023) use a simplified kinematic match methodology (matching the impact position at a single point) coupled with a deformable drop to model axisymmetric rebounds. They obtained good agreement with experiments and DNS, enabling a unified understanding of hte bouncing dynamics in their target inertio-capillary regime.
@@
We next present the derivation of the system, considering each fluid region independently in Section 2. We then couple the regions together and present a numerical approach for simulating the results in Section 3. Finally, we will use DNS techniques that have been previously verified against experiments to determine the accuracy of our model, and demonstrate that it can produce efficient simulations of complex multiple-rebound dynamics off a vibrating bath in Section 4 and Section 5, respectively.
2 Model
In this section we present a model to capture a deformable liquid droplet supported by a thin film of air which allows it to bounce off a deep liquid bath. Through considering the air, drop and bath separately, we will begin by assuming momentum and mass equations in each region, then include distinct approximations required for each region respectively, and describe the coupling through the pressure generated within the air layer from the thin film equation.
Consider a two-dimensional liquid droplet moving through an air-filled upper half plane towards a deep bath of infinite depth and width. We restrict the centre of mass of the droplet to vertical motion along the $z$-axis, introducing a symmetry to the problem with respect to the $z$ axis. Introducing the subscript notation $\cdot_{\alpha}$ to denote the different fluid regions, where $\alpha\in\{a,b,d\}$ are the air, bath, and droplet respectively, the velocity of each fluid is given by $\mathbf{u}_{\alpha}$, the pressure is given by $p_{\alpha}$, and $\mu_{\alpha}$, $\rho_{\alpha}$ and $g$ denote the dynamic viscosity, density and gravity, respectively.
@@ 2 Model
where $\sigma$ is the surface tension coefficient, $\kappa$ denotes the curvature of the free surface whose unit normal vector is given as $\mathbf{n}_{\beta}$, and $\tau=(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})$ is the stress tensor. In the far field for the air and the bath, the velocities vanish, and pressures tend to their hydrostatic balance values. We shall now consider the flows within the bath, air, and liquid domains separately, using the relevant approximation the respective domain requires.
2.1 Linear quasipotential bath model
Following methodology used in Milewski et al. (2015); Galeano-Rios et al. (2017, 2021) and Alventosa et al. (2023) to describe the waves generated by a bath-droplet interaction, we proceed by assuming a fluid bath with small free-surface perturbations from the equilibrium $z=0$ denoted by $z=\eta_{b}(x,t)$. As the liquid is incompressible we can take the velocity field of the bath as $\mathbf{u}_{b}=(u_{b},w_{b})=\nabla\phi+\nabla^{\perp}\psi$, where $\nabla\phi$ denotes the irrotational velocity component, and $\psi$ a small vortical correction arising from viscous effects. Here we have used the notation $\nabla^{\perp}=(-\partial_{z},\partial_{x})$. Due to the assumption that disturbances are small, the system (1) – (5) results in the following linearized problem
@@ 2.1 Linear quasipotential bath model
The arguments in Dias et al. (2008) (building on the viscous damping rates for water waves obtained by Lamb (1924)) involve a smallness assumption in the boundary layer flow represented by $\psi$ and a boundary layer scaling for $\psi$ in its vertical dependence. The boundary conditions then simplify to
@@ 2.2 Lubricating air layer model
When the droplet of undisturbed radius $R_{0}$ is sufficiently far from the bath, the two liquid regions have a negligible effect on each other. However, when the height between the drop and the bath is of $\mathcal{O}(\varepsilon R_{0})$, where $\varepsilon\ll 1$, they can start to affect each other and we are able to proceed by considering the air layer and the pressure within as a lubrication region (Hicks & Purvis, 2013).
@@ 2.3.2 Droplet deformation
The equations which govern the evolution of drop deformations follow the methodology first used in Lamb (1924), and more recently described in Aalilija et al. (2020) and Alventosa et al. (2023). We describe the motion of fluid within the droplet in polar coordinates, with origin at the centre of mass of the drop and $\theta=0$ in the negative $z$-direction. The free surface of the droplet at rest lies along $r=R_{0}$, with small perturbations to this given by $\eta_{d}(\theta,t)$.
@@
are viscous damping and frequency terms, respectively. We note that these correspond to the values in Aalilija et al. (2020) using the energy method of Lamb (1924), but were obtained directly from the quasi-potential approximation. The surface tangential velocity which appears in the lubrication equation (30) can be recovered by the leading order
@@ 3 Numerical Methods
In the previous section we introduced the equations that govern each of the three fluid regions and how they fit within a full lubrication-mediated quasi-potential fluid model. We shall now introduce the numerical methods used to evolve this system during this section. We will also describe a high-accuracy DNS implementation for the full system, which can be used as a benchmark for the various approximations in our system.
@@ 3.1 Solving the Lubrication-mediated quasi-potential model
In the numerical results we find that the lubrication layer thickness is insensitive to variation for a range of small values of $\varepsilon$, thus the value $\varepsilon=10^{-2}$ is used to set the outermost point in the lubrication layer according to the formula above. In Moore (2021) the nondimensional value $(\mu_{a}/\rho_{b}R_{0}W_{0})^{1/3}$ obtained from the horizontal momentum balance in the layer sets the vertical to horizontal scaling of the air layer. In our cases this value is ranged between $5\times 10^{-2}<(\mu_{a}/\rho_{b}R_{0}W_{0})^{1/3}<10^{-1}$.
@@
Numerical time-stepping used the Matlab ode45 adaptive Runge-Kutta scheme. We also used the basic Runge-Kutta $4^{\textrm{th}}$ order scheme which utilised fixed timesteps for some computations to verify the results through varying the timestep chosen until good agreement between models was met (which occurred for $\delta t=1\times{10}^{-9}$ s) The adaptive timestepper was efficient as a small time step is needed during certain phases of impact: the system is highly sensitive at first contact between the drop and bath, and when the lubrication layer begins to retract. However, longer time steps can be taken during the remainder of the rebound, and a much larger ones prior to, or after contact. We have found that the droplet deformation can create instabilities which result in unphysical contact, particularly at higher impact speeds and larger droplets. The instabilities can be avoided by adding artificial damping (through an increased viscosity) to the droplet equations during impact. Solid impacts did not suffer from this instability.
@@ 3.2 DNS
The DNS framework for this study has been designed using the Basilisk open-source code infrastructure (Popinet, 2009, 2015), using a finite volume discretisation scheme and second order adaptive in time and space capabilities which benefit the multi-scale nature of the problem. The interplay between inertial, capillary and viscous forces make the target problem especially challenging, with Basilisk a particularly suitable choice in this context in view of its versatile setup (Popinet, 2018). Variations of our implementation have been used successfully in the past within an axisymmetric formulation as part of studies of impacting superhydrophobic solids (Galeano-Rios et al., 2021) and, more recently, liquid drops (Alventosa et al., 2023), with the present implementation representing a two-dimensional counterpart constructed to enable reduced-dimensional model comparisons with the deformable droplet model proposed in Section 2. While previous multi-methodological investigations contained rigorous validation efforts, we have consolidated our confidence in the robustness of our code in the present configuration as well. Both trajectory metrics and detailed information such as gas film thickness, anticipated to be of $\mathcal{O}(1)\ \mu$m based on the recent work of Tang et al. (2019), have been monitored in view of reaching mesh independence to a suitable level of accuracy. This required the usage of a dedicated non-coalescence setup (Sanjay et al., 2023), which ensures no artificial merging of the drop and pool occurs owing to grid resolution limitations. This required the testing of minimum cell sizes down to sub-micron scales.
@@
The computational box is set up to measure $32R_{0}$ in each dimension, and is half-filled with liquid, leading to sufficiently large domains in terms of both depth and lateral extent to prevent bottom effects and outgoing wave reflections to affect the results of interest. Resolution level $12$, pertaining to a setting with $2^{12}$ cells per dimension for the minimum cell size, was found to be sufficient to ensure reliable results in our parameter regimes of interest following careful numerical experimentation. The adaptive refinement strategy focused gridnodes in regions with fluid-fluid interfaces, as well as regions with large changes in magnitude of the velocity components, ensuring sufficient resolution in regions around the drop, as well as near the impact region, and supporting the accurate capturing of the outgoing waves in the pool. This led to runs characterised by $\mathcal{O}(10^{5})$ gridnodes (a highly significant, and indeed necessary, decrease from the $\mathcal{O}(10^{7})$ uniform grid counterpart setup) and approximately $\mathcal{O}(10^{2})$ CPU hours per run, with the adaptive timestepping becoming particularly relevant in the delicate stages in which the drop and the pool are in contact.
@@
The above data-rich framework provided access to carefully resolved drop and bath interface coordinates, as well as other general information of interest such as velocities and pressure field data, which enables verification of key hypotheses and ensures the creation of a strong foundation for the deformable drop model comparisons, which are expanded upon in Section 4.

### v3

@@ Abstract
In order for a droplet to rebound rather than coalesce with a liquid bath, a layer of gas must persist throughout the impact. This gas, typically an air layer acts as a lubricant to the system and permits a pressure transfer between the two liquid bodies. Through considering separately the bath, air, and drop regions of fluid, we introduce a fully coupled reduced dynamic model of two-dimensional droplets (i.e. cylindrical geometry) rebounding off liquid baths, which incorporates an evolving lubricating air layer. Numerical solutions of the lubrication-mediated model are compared to dedicated direct numerical simulation of the Navier-Stokes equations. The reduced model captures rebounding dynamics well in the regime where it is most relevant: for low-speed impacts of small droplets, where capillary forces are important. Numerically, the reduced model is efficient, allowing for the computation of multiple rebounds and of long time dynamics of droplets rebounding on a vibrating bath. Furthermore, the lubrication-mediated model is able to provide detailed information within the air layer such as pressure and lubrication-layer geometry, which is usually omitted from reduced models.
@@ 1 Background and motivation
Since the work of Worthington [1, 2], we have endeavoured to classify the different behaviours that arise when droplets impact a liquid surface. An important question is to distinguish under which conditions a droplet will rebound versus coalesce or splash [3], especially in industrial processes such as spray injections [4].
@@
The ability for a droplet to rebound depends entirely on whether the layer of air that separates the two fluids remains unbroken throughout the impact. In the event that this air layer becomes sufficiently thin, the Van der Waals forces between the two liquid regions act, and the liquid pierces through the air layer and initiates coalescence [5]. For rebound to occur, it is a delicate balance between the system having enough restoring force to jettison the drop upwards after a cushioned impact, but not so much that the air layer is breached due to the inertia of the drop before the liquid surface has time to force the droplet upwards. "inviscid fluid model for the water entry of a rigid, solid body into a quiescent pool
@@
Industrial interest in solid impact dynamics were initially heavily motivated by military interest in sea-plane landings and slamming of boat hulls, leading to the development of Wagner theory [6, 7] as a foundational inviscid model for solid body entry into quiescent liquids. The role of the air layer in pre-impact dynamics of solid bodies on deep liquid baths was first developed by Smith et al. [8]. For this work we base ourselves on the investigation of Moore [9], who combined pre-impact cushioning with the Wagner model. We incorporate surface tension effects in drop oscillations and in wave generation and propagation in the bath, relevant to scenarios such as spray coating and ink-jet printing (see Derby[10]) where surface tension dominates small droplet behaviour.. Additionally, studies of droplet-droplet impacts have been motivated by the rich field of aerosols in a wide variety of natural processes from cloud seeding [11] to pathogen transfer through suspended droplets [12].
@@
This study is originally motivated by phenomena where a vertical vibration of the liquid bath permits sustained repeated droplet bouncing to occur [13]. These Faraday pilot-wave systems, named due to the Faraday waves generated at sufficiently strong vibration, can present a wide range of complex behaviours. To obtain a strongly coupled drop-bath system, the vibration is held below that which destabilises the free surface (the Faraday threshold). Within this regime, there are multiple drop-bath behaviours that can be experimentally achieved and modelled, and have been quantified in work such as those of Protiere et al. [14], and Milewski et al. [15]. As the forcing reaches near the Faraday threshold, the bouncing modes are destabilised, and through the asymmetry of impact a horizontal force is applied to the drop, permitting them to ‘walk’ along the surface of the bath, guided by the waves generated from previous bounces. This Faraday pilot-wave system has drawn parallels to the quantum theory described by de Broglie [16]. Dubbed the hydrodynamic quantum analogue, attempts have been made to recreate macro-scale analogies to quantum phenomena. For comprehensive reviews see the works of Bush [17, 18] and references therein. The nature of these studies has required long-time simulations of repeated bouncing, which quickly become computationally expensive. Efforts to minimise this cost include reducing the number of spatial dimensions, such as laid out in the work by Nachbin et al. [19], where a two-dimensional system was used to investigate the relationship between a droplet and the bath topography and the resultant tunnelling effects between bath cavities, or linking coupled droplet dynamics [20], and recent works have highlighted the importance of including vertical dynamics within these methods [21].
Two-dimensional configurations provide comparatively fast, detailed understanding of the key behaviours which are computationally expensive in three-dimensions. This approach can also be viewed as the ‘infinite-cylinder’ approximation of the geometry as a two dimensional droplet represents the cross-section of a cylinder. This consideration in itself can be well-motivated by physical phenomena such as water-walkers, whose long feet are cylindrical to leading order [22].
@@
The linear inviscid model for the oscillations of a droplet was developed in 1924 by Lamb[23]. The model has been validated experimentally by work such as in Becker et al. [24]. Considering droplets in zero gravity cases allows for the uninterrupted study of their oscillatory nature [25] and the longer viscous damping timescales. Comparisons between computations of the Navier-Stokes equations and linear theory in two- and three-dimensions can be found in Aalilija et al. [26]. Following Lamb, they used the inviscid solution and an energy argument to find linear damping rates. For bouncing droplets, Molác̆ek & Bush considered a simplified model where both bath and droplet deform equally as a logarithmic spring [27]. Terwagne et al. experimentally investigated the role of deformation in the droplet rebound dynamics when the drop and bath are in different regimes, and proposed a double mass-spring system to replicate this range of regimes in a drop-bath system [28]. While solid spheres can rebound in this regime [29], droplet deformation converts kinetic energy to drop deformations on contact with the bath, and then may revert it back to kinetic energy at take-off, increasing upwards propulsion.
@@
where here the Weber number indicates ratio of inertia to surface tension, the Reynolds number encapsulates the relative effect of inertia to viscosity, and the Ohnesorge number provides a relationship between viscosity and surface tension. There are several other parameters which could play a role; in particular for larger drops and longer waves the Bond number $Bo=\rho gR_{0}^{2}/\sigma$ gains importance. Even small changes to these parameters will affect the overall droplet behaviour, including whether the drops will bounce, or coalesce [30]. The work presented within this paper will focus around the low-Weber and the quasi-inviscid low-Ohnesorge limit, in order to capture smooth rebound dynamics with small disturbances [5].
The investigation of single rebounds in this regime has thrived over recent years, with the main goal being computationally efficient models that faithfully reproduce the experimental and DNS results. The ‘kinematic match’ method for axisymmetric [31] and non-axisymmetric impacts [32] modelled the droplet as a sphere, with the pressure on its contact surface obtained as a constraint for matching the shape of the sphere’s and the surface of the bath. This method captured details of bouncing and walking droplets, matching experiments, and was later used in comparisons of solid sphere impacts [29]. The model neglected droplet deformation and the air layer. Alventosa et al. [33] use a simplified kinematic match methodology (matching the impact position at a single point) coupled with a deformable drop to model axisymmetric rebounds. They obtained good agreement with experiments and DNS, enabling a unified understanding of the bouncing dynamics in their target inertio-capillary regime (vanishing influence of viscosity and gravitational forces described by sufficiently small Ohnesorge and Bond numbers, i.e. $\text{Oh}\ll 1$ and $\text{Bo}\ll 1$).
@@
We next present the derivation of the system, considering each fluid region independently in Section 2. We then couple the regions together and present a numerical approach for simulating the results in Section 3. Finally, we will use DNS techniques that have been previously verified against experiments to determine the accuracy of our model, and demonstrate that it can produce efficient simulations of complex multiple-rebound dynamics off a vibrating bath in Section 4 and Section 5, respectively. We conclude with a discussion of the model and consideration for future avenues of study in Section 6.
2 Mathematical model
Consider a two-dimensional liquid droplet, as though an infinite-cylinder in the out-of-plane direction, moving through an air-filled upper half plane towards a deep bath of infinite depth and width as shown in figure 1. We restrict the centre of mass of the droplet to vertical motion along the $z$-axis, introducing a symmetry to the problem with respect to the $z$ axis. Introducing the subscript notation $\cdot_{\alpha}$ to denote the different fluid regions, where $\alpha\in\{a,b,d\}$ are the air, bath, and droplet respectively, the velocity of each fluid is given by $\mathbf{u}_{\alpha}$, the pressure is given by $p_{\alpha}$, and $\mu_{\alpha}$, $\rho_{\alpha}$ and $g$ denote the dynamic viscosity, density and gravity, respectively.
@@ 2 Mathematical model
where $\sigma$ is the surface tension coefficient, $\kappa$ denotes the curvature of the free surface whose unit normal vector is given as $\mathbf{n}_{\beta}$, and $\tau=(\nabla\mathbf{u}+\nabla\mathbf{u}^{T})$ is the rate of strain tensor. In the far field for the air and the bath, the velocities vanish, and pressures tend to their hydrostatic balance values. We shall now consider the flows within the bath, air, and liquid domains separately, using the relevant approximation the respective domain requires.
2.1 Linear quasi-potential bath model
Following methodology well established in the literature [15, 31, 29, 33] to describe the waves generated by a bath-droplet interaction, we proceed by assuming a fluid bath with small free-surface perturbations from the equilibrium $z=0$ denoted by $z=\eta_{b}(x,t)$. As the liquid is incompressible we can take the velocity field of the bath as $\mathbf{u}_{b}=(u_{b},w_{b})=\nabla\phi+\nabla^{\perp}\psi$, where $\nabla\phi$ denotes the irrotational velocity component, and $\psi$ a small vortical correction arising from viscous effects. Here we have used the notation $\nabla^{\perp}=(-\partial_{z},\partial_{x})$. Due to the assumption that disturbances are small, the system (1) – (5) results in the following linearized problem
@@ 2.1 Linear quasi-potential bath model
The arguments in Dias et al [34] (building on the viscous damping rates for water waves obtained by Lamb [23]) involve a smallness assumption in the boundary layer flow represented by $\psi$ and a boundary layer scaling for $\psi$ in its vertical dependence. This is explained in more detail in Appendix B. The boundary conditions then simplify to
@@ 2.2 Lubricating air layer model
When the droplet of undisturbed radius $R_{0}$ is sufficiently far from the bath, the two liquid regions have a negligible effect on each other. However, when the height between the drop and the bath is of $\mathcal{O}(\varepsilon R_{0})$, where $\varepsilon\ll 1$, they can start to affect each other and we are able to proceed by considering the air layer and the pressure within as a lubrication region [35].
@@ 2.3.2 Droplet deformation
The equations which govern the evolution of drop deformations follow the methodology first used by Lamb [23], and more recently described by Aalilija et al [26] and Alventosa et al [33]. We describe the motion of fluid within the droplet in polar coordinates, with origin at the centre of mass of the drop and $\theta=0$ in the negative $z$-direction. The free surface of the droplet at rest lies along $r=R_{0}$, with small perturbations to this given by $\eta_{d}(\theta,t)$.
@@
are viscous damping and frequency terms, respectively. We note that these correspond to the values in Aalilija et al[26] using the energy method of Lamb[23], but were obtained directly from the quasi-potential approximation. The surface tangential velocity which appears in the lubrication equation (30) can be recovered by the leading order
@@ 3 Numerical methods
In the previous section we introduced the equations that govern each of the three fluid regions and how they fit within a full lubrication-mediated (LM) quasi-potential fluid model. We shall now introduce the numerical methods used to evolve this system during this section. We will also describe a high-accuracy direct numerical simulation (DNS) implementation for the full multi-fluid system, which can be used as a benchmark for the various approximations in our proposed methodology. Fully open-source codebases for the LM and DNS approaches presented in this section, as well as for the LM extension to Faraday systems detailed in Section 5, can be found in the dedicated GitHub repository: https://github.com/rcsc-group/BouncingDropletsLiquid2D [36].
@@ 3.1 Solving the lubrication-mediated quasi-potential model
In the numerical results we find that the lubrication layer thickness is insensitive to variation for a range of small values of $\varepsilon$, thus the value $\varepsilon=0.01$ is used to set the outermost point in the lubrication layer according to the formula above. In Moore [9] the nondimensional value $(\mu_{a}/\rho_{b}R_{0}W_{0})^{1/3}$ obtained from the horizontal momentum balance in the layer sets the vertical to horizontal scaling of the air layer. In our cases this value ranges between $0.10.05<(\mu_{a}/\rho_{b}R_{0}W_{0})^{1/3}<$ 0.1.
@@
Numerical time-stepping used the Matlab ode45 adaptive Runge-Kutta scheme. We also used the basic Runge-Kutta $4^{\textrm{th}}$ order scheme which utilised fixed timesteps for some computations to verify the results through varying the timestep chosen until good agreement between models was met (which occurred for $\delta t=10^{-9}$s) The adaptive timestepper was efficient as a small time step is needed during certain phases of impact: the system is highly sensitive at first contact between the drop and bath, and when the lubrication layer begins to retract. However, longer time steps can be taken during the remainder of the rebound, and a much larger ones prior to, or after contact. We have found that the droplet deformation can create instabilities which result in unphysical contact, particularly at higher impact speeds and larger droplets. The instabilities can be avoided by adding artificial damping (through an increased viscosity) to the droplet equations during impact. Solid impacts did not suffer from this instability.
@@ 3.2 Direct numerical simulation
The DNS framework for this study has been designed using the Basilisk open-source code infrastructure [37, 38], using a finite volume discretisation scheme and second order adaptive in time and space capabilities which benefit the multi-scale nature of the problem. The interplay between inertial, capillary and viscous forces make the target problem especially challenging, with Basilisk a particularly suitable choice in this context in view of its versatile setup [39]. Variations of our implementation have been used successfully in the past within an axisymmetric formulation as part of studies of impacting superhydrophobic solids [29] and, more recently, liquid drops [33], with the present implementation representing a two-dimensional counterpart constructed to enable reduced-dimensional model comparisons with the deformable droplet model proposed in Section 2. While previous multi-methodological investigations contained rigorous validation efforts, we have established the robustness of our code in the present configuration as well. Both trajectory metrics and detailed information such as gas film thickness, anticipated to be of $\mathcal{O}(1)\ \mu$m based on the recent experimental work of Tang et al [40], have been monitored in view of reaching mesh independence to a suitable level of accuracy. This required the usage of a dedicated non-coalescence setup [41], which ensures no artificial merging of the drop and pool occurs owing to grid resolution limitations. This required the testing of minimum cell sizes down to sub-micron scales.
@@
The computational box is set up to measure $32R_{0}$ in each dimension, and is half-filled with liquid, leading to sufficiently large domains in terms of both depth and lateral extent to prevent bottom effects and outgoing wave reflections to affect the results of interest. Resolution level $12$, pertaining to a setting with $2^{12}$ cells per dimension for the minimum cell size, was found to be sufficient to ensure reliable results in our parameter regimes of interest following careful numerical experimentation. The adaptive refinement strategy focused gridnodes in regions with fluid-fluid interfaces, as well as regions with large changes in magnitude of the velocity components, ensuring sufficient resolution in regions around the drop, as well as near the impact region, and supporting the accurate capturing of the outgoing waves in the pool. We note to have also made use of an interpolation method employing a harmonic mean formulation for the viscosity in the design of our implementation, as well as interface filtering techniques, owing to the relatively large contrast of the physical properties between the fluids used which may generate undesired instabilities, particularly in a rapidly evolving multi-scale structure such as the one present here. This led to runs characterised by $\mathcal{O}(10^{5})$ gridnodes (a highly significant, and indeed necessary, decrease from the $\mathcal{O}(10^{7})$ uniform grid counterpart setup) and approximately $\mathcal{O}(10^{2})$ CPU hours per run, with the adaptive timestepping becoming particularly relevant in the delicate stages in which the drop and the pool are in contact.
@@
The above data-rich framework provides access to carefully resolved drop and bath interface coordinates, as well as other general information of interest such as velocities and pressure field data, which enables verification of key hypotheses and ensures the creation of a strong foundation for the deformable drop model comparisons, which are expanded upon in Section 4.